A nozzle is a device that is used to accelerate a flowing fluid to higher velocity. If steam flows through an adiabatic nozzle at a steady state entering the nozzle with a pressure of 0.6 Mpa a velocity of 30 m/s and tempreture
1
Expert's answer
2021-07-22T10:02:46-0400

"\frac{pv^2}T=const,"

"\frac{p_1v_1^2}{T_1}=\frac{p_2v^2_2}{T_2},"

"v_2=v_1\sqrt{\frac{p_1T_2}{p_2T_1}}=30\cdot\sqrt{\frac{0.6\cdot 623}{0.2\cdot 673}}=50~\frac ms."